The length of the longest side of a triangle is twice the length of the shortest side. The length of the middle-sized side is five more than the length of the shortest side. Find the lengths of the sides of the triangle if the perimeter is 33 feet.
Answer by checkley77(12844)

LONGEST SIDE=2X
SHORTEST SIDE=X
MIDDLE SIDE=X+5
2X+X+X+5=33
4X=33-5
4X=28
X=28/4
X=7 FOR THE SHORTEST SIDE.
7+5=12 FOR THE MIDDLE SIDE
2*7=14 FOR THE LONGEST SIDE.
PROOF:
7+12=14=33
33=33